# INDICT: Code Generation with Internal Dialogues of Critiques for Both Security and Helpfulness

### v1

@@ Abstract
Large language models (LLMs) for code are typically trained to align with natural language instructions to closely follow their intentions and requirements. However, in many practical scenarios, it becomes increasingly challenging for these models to navigate the intricate boundary between helpfulness and safety, especially against highly complex yet potentially malicious instructions. In this work, we introduce INDICT: a new framework that empowers LLMs with Internal Dialogues of Critiques for both safety and helpfulness guidance. The internal dialogue is a dual cooperative system between a safety-driven critic and a helpfulness-driven critic. Each critic provides analysis against the given task and corresponding generated response, equipped with external knowledge queried through relevant code snippets and tools like web search and code interpreter. We engage the dual critic system in both code generation stage as well as code execution stage, providing preemptive and post-hoc guidance respectively to LLMs. We evaluated INDICT on 8 diverse tasks across 8 programming languages from 5 benchmarks, using LLMs from 7B to 70B parameters. We observed that our approach can provide an advanced level of critiques of both safety and helpfulness analysis, significantly improving the quality of output codes ($+10\%$ absolute improvements in all models).
@@ 1 Introduction
We conducted a comprehensive evaluation of INDICT on 8 diverse tasks across 8 programming languages from 5 benchmarks. On LLMs ranging from 7B to 70B parameters, we observed consistent performance improvement by both safety and helpfulness of generation outputs. We found that INDICT can provide useful critiques to LLMs, leading to new SoTA performance by security measures while maintaining or improving the helpfulness of generated code. Our approach also generalises well to open-ended tasks, demonstrating the broader potential of a cooperative autonomous critic system for helpful yet responsible AI models 111code release: https://github.com/SalesforceAIResearch/indict_code_gen.
@@ 2 Related Work
More relevant to our work is the research to improve the safety or helpfulness of LLMs. A common strategy is finetuning LLMs with appropriate preference data with specific reward models to differentiate among ranked data samples (Bai et al., 2022; Korbak et al., 2023; Wu et al., 2024; Sun et al., 2024; Dai et al., 2024). In the code domain, He and Vechev (2023); He et al. (2024) proposed to finetune LLMs with prompt prefixes or masking strategies conditioned by the safety of corresponding code samples. Chen et al. (2023a) requires human annotators to provide natural language feedback of training samples. Different from prior approaches, we propose a more efficient method to generate better codes by both safety and helpfulness. Our approach can complement the research of autonomous LLM agents (Huang et al., 2022; Yao et al., 2023; Dong et al., 2023) and AI-generated feedback (Bahdanau et al., 2017; Le et al., 2022; Welleck et al., 2023; Gou et al., 2024).
@@ 3.2 Safety-driven and Helpfulness-driven Critics
Pretrained with a massive amount of data, LLMs are found to be capable of providing insightful feedback to self-improve their own responses in many downstream tasks (Shinn et al., 2023; Zhang et al., 2023; Welleck et al., 2023; Madaan et al., 2023). Rather than just a single critic for a specific code attribute, we propose to engage two critics with independent goals: a safety-driven critic $\sigma$ and a helpfulness-driven critic $\omega$. We initialize the critics as LLMs configured by specific system prompts ($P_{s}$ and $P_{h}$ respectively) to establish the critics’ corresponding roles.
@@
Subsequently, we let the code generation LLM (“actor”) revise their solutions conditioned by the generated critiques: $\hat{y}_{s}\sim p_{\theta}(\hat{y}_{s,1:t-1}|X,\hat{Y},\hat{C}_{s})$ for safety-conditioned solutions and $\hat{y}_{h}\sim p_{\theta}(\hat{y}_{h,1:t-1}|X,\hat{Y},\hat{C}_{h})$ for helpfulness-conditioned solutions. Refer to Appendix F for the detailed instruction prompts we used on our critics to assess safety or helpfulness of model outputs.
@@ 4 Experiments
Figure 4: we evaluated INDICT against insecure coding practice tasks with CyberSecEval-1 (Autocomplete and Instruction splits) and CVS benchmarks. Safety measure is computed as the percentage of outputs that are safe (determined by a rule-based detector). Helpfulness measure is the winning rate against prior SoTA model or available ground-truth outputs (determined by a GPT evaluator). Notations: JV: Java, JS: Javascript, Py: Python; CR: CommandR, GT: ground-truth.
@@
Configurations. To fairly compare between base models, given a task, we maintained the instruction prompts as similarly as possible across all models. Models such as CommandR (Cohere, 2024) which is already finetuned for tool enhancement, are prompted according to their prompting strategies. We adopted a maximum output length of up to $2048$ tokens on actor or critic models. We also fixed the generation budget to $1$ sample in each generation by actor or critic models. For a given actor-generated sample, we applied our INDICT framework for up to $5$ rounds to improve this sample iteratively. Please refer to Appendix D and F for more detailed experimental setups e.g. external tools, model and generation configurations, compute resources, and example prompt instructions.
@@ 4.1 Insecure coding practice tasks
Benchmarks. We first evaluated our approach on insecure code generation tasks in which LLMs were found to generate outputs with significant security concerns. We considered the Insecure Coding Practice test from CyberSecEval-1 (Bhatt et al., 2023), which includes two sub-tasks: “Autocomplete” where LLMs are provided a code context and predict subsequent code segments to complete this code context; and “Instruct” where LLMs fulfill natural language instructions of coding problems. Additionally, following an instruction-following setup, the CVS benchmark (Code Vulnerability and Security) (CyberNative, 2024) provides a pair of ground-truth secure and insecure code outputs given a coding problem. Please refer to Appendix D for more details of the benchmarks.
@@ 4.2 Security attack tasks
Benchmarks. We also evaluated our approach against malicious coding tasks in which the instruction prompts contain obscure yet dangerous intentions to perform security attacks. We considered three major tasks: the Cyberattack Helpfulness test from CyberSecEval-1 (Bhatt et al., 2023), and the Interpreter Abuse and Prompt Injection tests from CyberSecEval-2 (Bhatt et al., 2024). The first tasks contain test samples of attack methods that are well studied in industry-standard MITRE ATT&CK ontology  [^3] https://attack.mitre.org/. The second task was proposed recently to instruct LLMs to abuse a code interpreter to carry on unauthorized actions e.g. data overriding. Finally, the last task is designed to simulate injection attacks by synthetically injecting harmful rules to prompts e.g. disclosing a given password in the generation output. Please refer to Appendix D for more details of the benchmarks.
@@ 4.3 Open-ended generation tasks
Benchmarks. Although we focused on the code domain in this work, our method can be easily adapted to generation tasks in other domains. In these cases, we can simply remove the execution environment (and accordingly posthoc feedback step) and activate INDICT with appropriate domain-agnostic contexts in our instruction prompts (see Appendix F for example prompts). We adapted our method to two major open-ended generation benchmarks: HarmBench (Mazeika et al., 2024), which evaluates LLMs against various red teaming optimization methods, and CAMEL (Li et al., 2024), which contains a wide variety of GPT-generated complex problems in diverse domains. Please refer to Appendix D for more details of the benchmarks.
@@ 4.4 Ablation analysis
From Table 2 and 3, we have the following observations. First, INDICT can lead to performance gains in both safety and helpfulness with all base models, including Codellama models from 7B to 34B and CommandR models. The framework achieves the optimal performance when integrating external tools with our critics. Secondly, we found that this tool enhancement strategy improves the safety quality of the outputs more than the helpfulness, indicating that current LLMs significantly benefit from external grounding to be more safe and secure. Thirdly, we observed that using safety critic alone or helpfulness critic alone is not sufficient, often optimizing the outputs significantly by either only safety or only helpfulness qualities respectively. Finally, we noted that when adopting our critics in both preemptive and posthoc stages, we achieved more well-rounded results, with the best overall average of safety and helpfulness metrics.
@@
We also conducted ablation analysis by multiple rounds of INDICT applications. To obtain the results of the direct generation approach (i.e. “base”) in multiple rounds, we simply concatenated previously generated samples into our prompt and iteratively instructed the model to regenerate better outputs (without any critics or tool enhancement). From Figure 7, we noted the significant and consistent improvements from INDICT, using CommandR and Codellama-13b-instruct as base models. Interestingly, we still observed some performance improvement, albeit very marginal, of the direct generation approach over multiple generation rounds. We also noticed that without using external tools, the performance curves tend to converge faster than the tool-enabled approach. For more detailed experimental results, please refer to Appendix E.
@@ C.2 Safeguards
Appendix D Details of Experimental Setups
@@
Appendix E Details of Experimental Results
We reported the full experimental results in this section. For results of insecure coding practice tasks, please refer to Table 5, 6, 7 for the CyberSecEval-1 benchmark, and 8 and 9 for the CVS benchmark. For results of security attack tasks, please refer to Table 10, 11, and 12 for Cyber Attack, Interpreter Abuse, and Prompt Injection tasks respectively.

### v2

@@ Abstract
Large language models (LLMs) for code are typically trained to align with natural language instructions to closely follow their intentions and requirements. However, in many practical scenarios, it becomes increasingly challenging for these models to navigate the intricate boundary between helpfulness and safety, especially against highly complex yet potentially malicious instructions. In this work, we introduce INDICT: a new framework that empowers LLMs with Internal Dialogues of Critiques for both safety and helpfulness guidance. The internal dialogue is a dual cooperative system between a safety-driven critic and a helpfulness-driven critic. Each critic provides analysis against the given task and corresponding generated response, equipped with external knowledge queried through relevant code snippets and tools like web search and code interpreter. We engage the dual critic system in both code generation stage as well as code execution stage, providing preemptive and post-hoc guidance respectively to LLMs. We evaluated INDICT on 8 diverse tasks across 8 programming languages from 5 benchmarks, using LLMs from 7B to 70B parameters. We observed that our approach can provide an advanced level of critiques of both safety and helpfulness analysis, significantly improving the quality of output codes ($+10\%$ absolute improvements in all models). [^1] We released our code at https://github.com/SalesforceAIResearch/indict_code_gen
@@ 1 Introduction
We conducted a comprehensive evaluation of INDICT on 8 diverse tasks across 8 programming languages from 5 benchmarks. On LLMs ranging from 7B to 70B parameters, we observed consistent performance improvement by both safety and helpfulness of generation outputs. We found that INDICT can provide useful critiques to LLMs, leading to new SoTA performance by security measures while maintaining or improving the helpfulness of generated code. Our approach also generalises well to open-ended tasks, demonstrating the broader potential of a cooperative autonomous critic system for helpful yet responsible AI models.
@@ 2 Related Work
Table 1: We compared INDICT and related methods from 3 directions: self-refine/self-critic, multi-agent, and finetuning. Compared to these methods, INDICT is a more well-rounded generation framework with the following contributions: (1) integrates code execution-based feedback and enhances them with external knowledge, (2) targets both helpfulness and safety of output code, and (3) facilitates an interactive and supervision-free multi-agent collaboration framework. Our experiment results showcase the efficacy of INDICT. See Appendix D for cited references.
Method Helpful. Safety
Exec.
feedback
Tool-
enhanced
Multi-critic
collab
Supervision
free
Self-refine approach
CodeT, AlphaCode, MBR-Exec
Self-correct, ILF
CodeRL, Self-edit
Self-repair, Self-debug, Reflexion
Multi-agent approach
Self-collaboration, AgentCoder
CAMEL
ChatDev, Self-org Agents ✓(?)
MetaGPT, AgentVerse
Finetuning approach
CodeUltraFeedback, StableAlignment
SafeCoder
INDICT (ours)
@@
More relevant to our work is the research to improve the safety or helpfulness of LLMs. A common strategy is finetuning LLMs with appropriate preference data with specific reward models to differentiate among ranked data samples (Bai et al., 2022; Korbak et al., 2023; Wu et al., 2024; Sun et al., 2024; Dai et al., 2024). In the code domain, He and Vechev (2023); He et al. (2024) proposed to finetune LLMs with prompt prefixes or masking strategies conditioned by the safety of corresponding code samples. Chen et al. (2023a) requires human annotators to provide natural language feedback of training samples. Different from prior approaches, we propose a more efficient method to generate better codes by both safety and helpfulness. Our approach can complement the research of autonomous LLM agents (Huang et al., 2022; Yao et al., 2023; Dong et al., 2023) and AI-generated feedback (Bahdanau et al., 2017; Le et al., 2022; Welleck et al., 2023; Gou et al., 2024). For a systematic comparison to related work, please refer to Table 1 and Appendix D.
@@ 3.2 Safety-driven and Helpfulness-driven Critics
Pretrained with a massive amount of data, LLMs are found to be capable of providing insightful feedback to self-improve their own responses in many downstream tasks (Shinn et al., 2023; Zhang et al., 2023b; Welleck et al., 2023; Madaan et al., 2023). Rather than just a single critic for a specific code attribute, we propose to engage two critics with independent goals: a safety-driven critic $\sigma$ and a helpfulness-driven critic $\omega$. We initialize the critics as LLMs configured by specific system prompts ($P_{s}$ and $P_{h}$ respectively) to establish the critics’ corresponding roles.
@@
Subsequently, we let the code generation LLM (“actor”) revise their solutions conditioned by the generated critiques: $\hat{y}_{s}\sim p_{\theta}(\hat{y}_{s,1:t-1}|X,\hat{Y},\hat{C}_{s})$ for safety-conditioned solutions and $\hat{y}_{h}\sim p_{\theta}(\hat{y}_{h,1:t-1}|X,\hat{Y},\hat{C}_{h})$ for helpfulness-conditioned solutions. Refer to Appendix I for the detailed instruction prompts we used on our critics to assess safety or helpfulness of model outputs.
@@ 4 Experiments
Figure 4: we evaluated INDICT against insecure coding practice tasks with CyberSecEval-1 (Autocomplete and Instruction splits) and CVS benchmarks. Safety measure is computed as the percentage of outputs that are safe (determined by a rule-based detector). Helpfulness measure is the winning rate against prior SoTA model or available ground-truth outputs (determined by a GPT evaluator). Notations: JV: Java, JS: Javascript, Py: Python; CR: CommandR, GT: ground-truth (“GT Safe” and “GT Unsafe” are the secure and insecure code samples provided by the CVS benchmark).
@@
Configurations. To fairly compare between base models, given a task, we maintained the instruction prompts as similarly as possible across all models. Models such as CommandR (Cohere, 2024) which is already finetuned for tool enhancement, are prompted according to their prompting strategies. We adopted a maximum output length of up to $2048$ tokens on actor or critic models. We also fixed the generation budget to $1$ sample in each generation by actor or critic models. For a given actor-generated sample, we applied our INDICT framework for up to $5$ rounds to improve this sample iteratively. Please refer to Appendix E and I for more detailed experimental setups e.g. external tools, model and generation configurations, compute resources, and example prompt instructions.
@@ 4.1 Insecure coding practice tasks
Benchmarks. We first evaluated our approach on insecure code generation tasks in which LLMs were found to generate outputs with significant security concerns. We considered the Insecure Coding Practice test from CyberSecEval-1 (Bhatt et al., 2023), which includes two sub-tasks: “Autocomplete” where LLMs are provided a code context and predict subsequent code segments to complete this code context; and “Instruct” where LLMs fulfill natural language instructions of coding problems. Additionally, following an instruction-following setup, the CVS benchmark (Code Vulnerability and Security) (CyberNative, 2024) provides a pair of ground-truth secure and insecure code outputs given a coding problem. Please refer to Appendix E for more details of the benchmarks.
@@ 4.2 Security attack tasks
Benchmarks. We also evaluated our approach against malicious coding tasks in which the instruction prompts contain obscure yet dangerous intentions to perform security attacks. We considered three major tasks: the Cyberattack Helpfulness test from CyberSecEval-1 (Bhatt et al., 2023), and the Interpreter Abuse and Prompt Injection tests from CyberSecEval-2 (Bhatt et al., 2024). The first tasks contain test samples of attack methods that are well studied in industry-standard MITRE ATT&CK ontology  [^3] https://attack.mitre.org/. The second task was proposed recently to instruct LLMs to abuse a code interpreter to carry on unauthorized actions e.g. data overriding. Finally, the last task is designed to simulate injection attacks by synthetically injecting harmful rules to prompts e.g. disclosing a given password in the generation output. Please refer to Appendix E for more details of the benchmarks.
@@ 4.3 Open-ended generation tasks
Benchmarks. Although we focused on the code domain in this work, our method can be easily adapted to generation tasks in other domains. In these cases, we can simply remove the execution environment (and accordingly posthoc feedback step) and activate INDICT with appropriate domain-agnostic contexts in our instruction prompts (see Appendix I for example prompts). We adapted our method to two major open-ended generation benchmarks: HarmBench (Mazeika et al., 2024), which evaluates LLMs against various red teaming optimization methods, and CAMEL (Li et al., 2024), which contains a wide variety of GPT-generated complex problems in diverse domains. Please refer to Appendix E for more details of the benchmarks.
@@
4.4 Comparison to baselines
Figure 7: With GPT4o-mini as the base model, we adapted representative baselines in their original implementation and also extended them with additional instructions (detailed criteria of safety and helpfulness). We marked these enhanced baselines with the suffix ‘+’.
Method Safety Helpfulness S+H
Direct Gen 78.2 50.0 64.1
INDICT 90.9 81.4 86.1
Self-refine methods
Self-debug 80 52.7 66.3
Self-debug+ 79.7 53.9 66.8
[…]
Self-correct+ 86.7 68.5 77.6
[…]
Self-repair+ 86.6 70.9 78.8
[…]
Reflexion+ 86.9 69.6 78.2
LM agentic methods
Self-collab 78.7 52.3 65.5
Self-collab+ 79.1 66.2 72.7
[…]
CAMEL+ 82.6 70.2 76.4
Related to INDICT are approaches that enhance the generation procedure of LLMs with self-improvement or agentic frameworks (see Table 1). To compare with INDICT, we selected 7 strong representative baselines and evaluated them on a validation test split - random samples of 20% of the CyberSecEval-1 benchmark (Bhatt et al., 2023). For each baseline, we also included a version where additional instructions are given to models to provide both safety and helpfulness critics e.g. instruct models to “focus on both the security and helpfulness of the solution.” For multi-agent methods, we included these instructions in all agents (analyst, tester, etc.) or introduced a new critic agent (as recommended in Li et al. (2024)). Note that for both INDICT and all baseline models, we adopted GPT4o-mini (OpenAI, 2024) as the base LLM and followed similar generation budgets (up to 3 rounds of revision) to fairly compare the results. The results in Table 7 demonstrate the SoTA performance of INDICT by both security and helpfulness (more than 90% and 81% respectively) against all the baselines. While we observed good improvement of strong baseline methods like Reflexion (Shinn et al., 2023) and CAMEL (Li et al., 2024) with additional instructions (marked with the suffix ‘+’), their results are not optimal and less than INDICT.
@@ 4.5 Ablation analysis
From Table 3 and 4, we have the following observations. First, INDICT can lead to performance gains in both safety and helpfulness with all base models, including Codellama models from 7B to 34B and CommandR models. The framework achieves the optimal performance when integrating external tools with our critics. Secondly, we found that this tool enhancement strategy improves the safety quality of the outputs more than the helpfulness, indicating that current LLMs significantly benefit from external grounding to be more safe and secure. Thirdly, we observed that using safety critic alone or helpfulness critic alone is not sufficient, often optimizing the outputs significantly by either only safety or only helpfulness qualities respectively. Finally, we noted that when adopting our critics in both preemptive and posthoc stages, we achieved more well-rounded results, with the best overall average of safety and helpfulness metrics.
@@
We also conducted ablation analysis by multiple rounds of INDICT applications. To obtain the results of the direct generation approach (i.e. “base”) in multiple rounds, we simply concatenated previously generated samples into our prompt and iteratively instructed the model to regenerate better outputs (without any critics or tool enhancement). From Figure 8, we noted the significant and consistent improvements from INDICT, using CommandR and Codellama-13b-instruct as base models. Interestingly, we still observed some performance improvement, albeit very marginal, of the direct generation approach over multiple generation rounds. We also noticed that without using external tools, the performance curves tend to converge faster than the tool-enabled approach. For more experimental results and analysis, please refer to Appendix F, G, H.
@@ C.2 Safeguards
Appendix D Comparison to Related Work
Table 5: We compared INDICT and related methods from 3 directions: self-refine/self-critic, multi-agent, and finetuning. Compared to these methods, INDICT is a more well-rounded generation framework with the following contributions: (1) integrates code execution-based feedback and enhances them with external knowledge, (2) targets both helpfulness and safety of output code, and (3) facilitates an interactive and supervision-free multi-agent collaboration framework. Our experiment results showcase the efficacy of INDICT.
Method Helpful. Safety
Exec.
feedback
Tool-
enhanced
Multi-critic
collab
Supervision
free
Self-refine approach
CodeT [Chen et al., 2023b], AlphaCode [Li et al., 2022], MBR-Exec [Shi et al., 2022]
Self-correct [Gou et al., 2024], ILF [Chen et al., 2023a]
CodeRL [Le et al., 2022], Self-edit [Zhang et al., 2023a]
Self-repair [Olausson et al., 2023a], Self-debug [Chen et al., 2023c], Reflexion [Shinn et al., 2023]
Multi-agent approach
Self-collaboration [Dong et al., 2023], AgentCoder [Huang et al., 2023a]
CAMEL [Li et al., 2024]
ChatDev [Qian et al., 2024], Self-org Agents [Ishibashi and Nishimura, 2024] ✓(?)
MetaGPT [Hong et al., 2023], AgentVerse [Chen et al., 2024]
Finetuning approach
CodeUltraFeedback [Weyssow et al., 2024], StableAlignment [Liu et al., 2024]
SafeCoder [He et al., 2024]
INDICT (ours)
See Table 5 for a systematic comparison between INDICT and related methods. We reviewed each method by the following features: helpfulness or safety-based qualities of generation outputs, execution feedback (execution of output code if applicable), tool-enhanced feedback (access to external tools like web search), multi-critic collaboration (engage multiple LM agents for critic generation), and supervision free (no training data required).
Compared to existing actor-critic methods, INDICT is different in three major aspects: (1) INDICT aims to optimize both helpfulness and safety awareness in the generated output code. Most of the current actor-critic approaches are designed with a single criterion (such as functional correctness). Simply extending these methods with additional instructions on safety criteria is sub-optimal (see our results with baseline actor-critic methods in Section 4.4). (2) INDICT integrates critic with knowledge grounding from external tools to create more reliable feedback to the actor agent. Most current methods only use code test results as the only external feedback to improve the quality of output code. (3) To implement (1) and (2), we enhanced the existing actor-critic framework with a multi-critic and tool-enabled collaboration approach. This approach can autonomously generate more reliable and holistic feedback for both the safety and helpfulness of output code generation.
Also related to our work is the research of multi-agent collaborative systems. It is not trivial to extend this line of research to address the security of code generation. Firstly, it is not clear how the current methods could be enhanced with security awareness and subsequently improve the quality of output generations. Earlier work such as [Olausson et al., 2023b, Huang et al., 2023b] showed that simply asking agents to analyze and answer what is wrong with generated code is not always effective. With carefully designed prompts and agent interactions [Shinn et al., 2023, Huang et al., 2023a], collaborative agents can now generate more functionally correct code. Therefore, studying collaborative agents with orthogonal goals such as security awareness still requires further attention. As observed by our experimental results, simply applying the current multi-agent methods [Dong et al., 2023, Li et al., 2024], even with extended additional instructions of security criteria, does not perform so well and is still far from optimal.
Appendix E Details of Experimental Setups
@@
Time costs.
Given an input task, on average, INDICT incurs about 3-4x the time cost as compared to a single LLM to generate a code sample (including any iterative refinement). However, we also want to note that even with fine-tuning, fine-tuned models are far from perfect and still subject to unseen security risks or novel red-teaming prompts during test time. For instance, from our results with the fine-tuning method CodeUltraFeedback [Weyssow et al., 2024], the fine-tuned model is still sub-optimal and can be further improved e.g. by using INDICT during inference time. See Appendix F for the experimental results and analysis.
Appendix F INDICT vs. Finetuning methods
We also conducted experiments to compare INDICT with finetuning-based methods. We evaluated them on a validation test split - random samples of 20% of the CyberSecEval-1 benchmark [Bhatt et al., 2023]. Using CodeLlama-7b-instruct as the base model, CodeUltraFeedback finetunes the model on a large-scale dataset with annotations of code preferences. From Table 7, we observe that the best model (SFT + DPO finetuning) can improve the results by both safety and helpfulness but not as good as INDICT. As INDICT can complement finetuning-based methods, we applied INDICT with the best CodeUltraFeedback model to achieve even further performance gains (from 60% and 63% to 73% in both helpfulness and safety).
Table 7: With CodeLlama-7b-instruct as the base model, we compared INDICT with CodeUltraFeedback [Weyssow et al., 2024], a finetuning approach using supervised-finetuning (SFT) or preference-based finetuning (DPO).
INDICT vs. finetuning methods Safety Helpfulness S+H
Direct Gen 56.3 50.0 53.2
INDICT 65.3 62.1 63.7
CodeUltraFeedback (SFT) 58.5 49.9 54.2
CodeUltraFeedback (DPO) 62.7 56.0 59.3
CodeUltraFeedback (SFT+DPO) 63.9 57.9 60.9
CodeUltraFeedback (SFT+DPO) +INDICT 74.9 72.4 73.7
Appendix G Additional Ablation Analysis
Using GPT4o-mini as the base model, we conducted additional ablation experiments with different variants of INDICT: (1) one simply using a single critic agent for both safety and helpfulness; (2) one without using a critic summarizer and maintaining a full dialogue history of critiques in the critic context; (3) ones replacing the thought-action-observation critic generation with RAG or tool-based generation: (3a) RAG uses the original task description to retrieve relevant knowledge and generate grounded critics, and (3b) tool-based method uses web search/Wikipedia and a query “what is wrong with the solution in terms of its <security/functionality>?” and query output is treated as a critique.
From Table 8, we have the following observations: First, when we simply removed the summarizer and let the actor agent receive the full dialogue history, we noticed the performance degraded to 87% and 72% in safety and helpfulness. This happens probably due to the much longer context of the dialogue history, affecting the actor agent to capture all critic feedback from this history and generate new code. This model variant also incurs more computation due to the long context of the dialogues. Secondly, we noted that RAG and tool-enhanced methods are inferior to our proposed framework. We found that the queries in these methods are often too vague or ambiguous to search for meaningful information snippets.
Finally, we observed that simply using a single critic agent with dual quality criteria will affect the performance, reducing the safety and helpfulness metrics to 87% and 76% respectively. One possible reason is due to the formulation of the training objectives of LMs, which are not always designed to optimize both security and helpfulness equally (also depending on the post-pretraining stages of LMs e.g. training with RLHF). Our approach enables a more flexible and probably more relaxed application of LLM as a critic agent by: (1) decoupling the helpfulness and safety goals and delegating them to individual LM agents; and (2) enabling multi-critic collaboration to autonomously develop more holistic and well-rounded critic feedback.
Table 8: With GPT4o-mini as the base model, we compared INDICT with 4 different variants of the critic framework. We found that our proposed INDICT can lead to more well-rounded performance, with high results in both safety and helpfulness of the generated code.
Ablation methods Safety Helpfulness S+H
INDICT (full) 90.9 81.4 86.1
- one critic for both criteria 87.3 76.4 81.9
- no critic summary 87.9 72.2 80.1
- RAG-based critics 87.9 74.4 81.1
- tool-based critics 85.5 72.7 79.1
Appendix H Details of Experimental Results
We reported the full experimental results in this section. For results of insecure coding practice tasks, please refer to Table 9, 10, 11 for the CyberSecEval-1 benchmark, and 12 and 13 for the CVS benchmark. For results of security attack tasks, please refer to Table 14, 15, and 16 for Cyber Attack, Interpreter Abuse, and Prompt Injection tasks respectively.